# On metric choice in dimension reduction for Fréchet regression

### v1

@@ Abstract
Fréchet regression is becoming a mainstay in modern data analysis for analyzing non-traditional data types belonging to general metric spaces. This novel regression method utilizes the pairwise distances between the random objects, which makes the choice of metric crucial in the estimation. In this paper, the effect of metric choice on the estimation of the dimension reduction subspace for the regression between random responses and Euclidean predictors is investigated. Extensive numerical studies illustrate how different metrics affect the central and central mean space estimates for regression involving responses belonging to some popular metric spaces versus Euclidean predictors. An analysis of the distributions of glycaemia based on continuous glucose monitoring data demonstrate how metric choice can influence findings in real applications.
keywords:
Keywords: $L_{p}$ spaces, distributional distances, network distances, diabetes mellitus, continuous glucose monitoring
@@ 1 Introduction
Because $Y$ is allowed to belong to a general metric space, including those that do not support vector space properties, Fréchet regression is applicable in many fields. Some applications of this novel regression can be found in the study of mortality profiles in Petersen & Müller (2019), the influence of age on the development of brain myelination in Petersen et al. (2019), and the association between age, height, and weight on the structural connectivity of the human brain in Soale & Dong (2023), to name a few.
@@
The rest of the paper is organized as follows. In Section 2, we provide a brief background on metric and pseudometric spaces. We also describe different metric choices for various spaces. In Section 3, we give a general background to Fréchet sufficient dimension reduction, followed by an extensive simulation studies in Section 4. We present a real application in Section 5 and conclude the paper in Section 6 with some discussion. All other technical details are relegated to the Appendix.
@@ 2.1 General background
Furthermore, it is important to note that regardless of whether we are dealing with a metric or a pseudometric space, different metric functions in a given space can have dramatic effects on the geometry of the space. For instance, in the normed space, we use the $\ell_{p}$ norm, denoted as .$\lVert.\rVert_{p}$, where $p\geq 1$ is some real number. In this space, the choice of $p$ determines the geometry as shown in Figure 1. The Figure also reveals some “nesting” or “ordering” behavior of the $\ell_{p}$ norm.
@@
Proposition 2.2.
@@
By Proposition 2.2, $\ell_{2}$ is the most restrictive of the $\ell_{p}$ norms, and in a finite normed space, the richness of $\mathbb{L}_{p}$ depends on the choice of metric. In particular, for $1\leq p_{1}<p_{2}\leq 2,\ \mathbb{L}_{p_{2}}\subset\mathbb{L}_{p_{1}}$ [Matoušek (2013)]. This containment relation can have significant effects on the central space estimation as we will show later. There are several other metrics in different metric spaces that are based on the $\ell_{p}$ norm.
2.2 Popular metric functions in some common metric spaces
Euclidean space
In this space, the response set $\bm{\Omega}_{Y}\in\mathbb{R}^{r}$, where $r\geq 1$ is some positive integer. The most common metric in this space is the $\ell_{p}$ norm of which we will focus on the $\ell_{1}$ and $\ell_{2}$ norms. By Proposition 2.2, in the finite Euclidean space, $\mathbb{L}_{2}(Y)\subset\mathbb{L}_{1}(Y)$. Besides these two metrics, for $r\geq 2$, we will also consider the Mahalanobis distance, which is a pseudometric that takes into account the dependence between the responses. The Mahalanobis distance is given by
@@ 2.2 Popular metric functions in some common metric spaces
Space of probability distributions
@@
By Proposition 2.4, when there is 1st-order stochastic dominance, $W_{1}(.)$ simply reduces to the difference in means. For example, for any two log-normal distributions $F_{1}(\mu_{1},\sigma)$ and $F_{2}(\mu_{2},\sigma)$, $F_{1}$ dominates $F_{2}$ if $\mu_{1}>\mu_{2}$. In such scenarios, using the $W_{1}(.)$ is preferable to $W_{2}(.)$ as the latter will likely introduce some noise in finite samples based on the decomposition in (4).
@@
Space of symmetric positive definite matrices
@@
[…]
@@ 3.1 A brief background
Definition 3.1 (Metric embedding).
@@
Theorem 3.2 emphasizes the importance of choosing the appropriate metric as it relates to information about the central space such as dimensionality, unbiasedness, Fisher consistency, and exhaustiveness. See Li (2018) for more details on these properties.
Theorem 3.3.
@@
For Theorem 3.3 to be satisfied, it suffices for the response embedding to be Lipschitz, which may not hold for all metrics defined on the same space. For instance, the Wasserstein distance between two normal distributions is Lipschitz but the Hellinger distance is not. Some metrics may also satisfy the Lipschitz condition locally. Moreover, for metrics without upper bounds, the embedding on scaled inputs is more likely to concentrate around the mean than the embedding based on the actual values. For instance, we expect the Frobenius norm of the difference between logged positive definite matrices to concentrate around the mean faster than the Frobenius norm between the actual matrices. Thus, it is important to evaluate metrics on case-by-case basis. The central mean space estimators are especially powerful when the concentration of the metric is Gaussian or approximately Gaussian.
@@ 4.1 Euclidean responses
Model $(n,p)$ Metric F-OLS sa-OLS FSIR sa-SIR
@@ 4.2 Space of univariate distributions
Here, all three models III, IV, and V are single-index with $\eta=\beta_{1}$. From Table LABEL:tab:Dist_sims, the Hellinger distance shows significant improvement in the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates for both models III and V over the Wasserstein metrics. For model III, the Wasserstein distances are not globally Lipschitz but may be locally Lipschitz when $\beta_{1}^{\top}{\bf X}<0$. Thus, we do not expect $W_{1}$ and $W_{2}$ to concentrate around the mean by Theorem 3.3. On the other hand, the Hellinger distance concentrates around the center. Moreover, the Hellinger distance is better suited for capturing the tail variations compared to the Wasserstein metrics. This is one of the reasons for the difference we see for the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates for model III. However, we see less difference in the ${\mathcal{S}}_{Y|{\bf X}}$ estimates as the central space is not restricted to the mean.
In model IV, the Hellinger distance performs poorly compared to Wasserstein metrics in both the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and ${\mathcal{S}}_{Y|{\bf X}}$ estimates. The 1-Wassertein performs slightly better than the 2-Wasserstein because any pair of log-normal distributions with same scale parameters exhibit stochastic dominance. Thus, based on Proposition 2.4, $W_{1}(.)$ captures just the difference in the means while $W_{2}(.)$ may introduce some noise from the difference in sample scale parameters and shape. The log-normal distributions are more symmetric, which makes the Wasserstein a better metric for capturing location differences than the Hellinger distance.
@@
Model $(n,p)$ Metric F-OLS sa-OLS FSIR sa-SIR
@@ 4.3 Space of positive definite matrices
Model $(n,p)$ Metric F-OLS sa-OLS FSIR sa-SIR
@@ 4.4 Space of networks
Model $(n,p)$ Metric F-OLS sa-OLS FSIR sa-SIR
@@
5 Glucose monitoring data
@@
Type 2 diabetes mellitus (T2DM) disease has become rampant across the world in recent decades. Many reasons have been attributed to this with one of the most prominent being the rise in obesity. Classifying a patient as diabetic is typically based on some threshold of the level of glucose in the blood (glycaemia), which could potentially lead to incorrect diagnosis. However, with recent advances in continuous glucose monitoring systems, we could improve the metrics for diagnosing type 2 diabetes mellitus. The goal of this study is to analyze the distribution of glycaemia and determine the factors influencing the distribution. This way, we take into account all aspects of the distribution including the center and tails.
@@
Our data comes from the continuous glucose monitoring of 208 patients selected from the University Hospital of Móstoles, in Madrid, during an outpatient visit from January 2012 to May 2015. Our response of interest is the distribution of glycaemia recorded every 5 seconds for each patient over at least a 24 hour period. Patients were followed until either the diagnosis of T2DM or end of study. For each patient, clinical variables including gender (male/female), age (years), body mass index (BMI; $kg/m^{2}$), basal glycaemia ($mg/dL$), haemoglobin A1c (HbA1c; %), and follow-up (days), were also recorded. We take these clinical variables as the predictors. Further details about the collection procedure can be found in Colás et al. (2019).
Next, we proceed to estimate the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and the ${\mathcal{S}}_{Y|{\bf X}}$ for the regression between the glycaemia distributions and the clinical variables. Since, we are dealing with distributional responses, we use the Wasserstein and Hellinger distances to see how they affect the estimated results. To compare the performance of the methods and metrics, we estimate the leave-one-out average prediction error for the conditional Fréchet mean in model (1).
@@
Let $\hat{\beta}$ denote the basis estimate. The average leave-one-out prediction error is given by
@@
where $\hat{m}_{\oplus}(\bm{x})$ is the sample estimate of $m_{\oplus}(\bm{x})$ and $\bm{x}^{(-i)}$ is $\bm{x}$ without the $i$th observation. The estimated bases for ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and ${\mathcal{S}}_{Y|{\bf X}}$ are provided in Tables 5 and 6, respectively.
Table 5: estimated bases of ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ based on different metrics
@@
Age -0.4812 -0.4674 -0.5620 -0.5509 -0.1832 -0.2291
BMI 0.0530 0.0665 -0.0469 -0.0350 -0.1245 -0.2994
Basal glycaemia -0.2185 -0.2337 -0.2257 -0.2408 0.4399 0.4702
HbA1c -0.4168 -0.4239 -0.3874 -0.3918 -0.2083 -0.2173
Follow-up 0.6536 0.6486 0.586 0.5831 -0.8092 -0.7657
$\delta$ 12.0313 12.0354 11.9249 11.9721 12.0558 12.0076
Based on the results in Table 5, the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates based on the 2-Wasserstein yield the smallest prediction error compared to the estimates based on the other metrics. Although their basis estimates appear to differ significantly for BMI, the prediction error based on 1-Wasserstein and Hellinger distance appear to be similar. Based on the FOLS($W_{2}$) estimate, which is the best, the subject’s age and follow-ups are the most influential in predicting the average glycaemia level, followed by the Hemoglobic A1C. BMI has the least influence on the average concentration of glucose in the blood. This result goes against the conventional knowledge about the relationship between BMI and glycaemia such as Hassan et al. (2021). However, this is not surprising at the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ focuses on only the center of the distributions. Therefore, we proceed to estimate the central space for broader relationships.
Table 6: estimated bases of ${\mathcal{S}}_{Y|{\bf X}}$ based on different metrics
@@
Age 0.6132 0.6142 -0.6951 0.6592 -0.2035 -0.2465
BMI 0.6440 0.6324 -0.6609 0.6768 0.0031 -0.0151
Basal glycaemia -0.5329 -0.6131 0.4556 -0.5093 0.5506 0.5063
HbA1c 0.3813 0.3246 -0.2946 0.2782 -0.0642 -0.1145
Follow-up -0.1794 -0.1576 0.1395 -0.1643 -0.7431 -0.7631
$\delta$ 12.0549 12.1245 12.0636 12.0620 12.1252 12.1392
Looking at the ${\mathcal{S}}_{Y|{\bf X}}$ estimates in Table 6, the 1-Wasserstein based on FSIR and the estimates based on the 2-Wasserstein appear to be better at predicting the glycaemia distribution than the rest. For the broader relationship between the distributions and the predictors, we see that BMI, age, and the basal glycaemia have the most influence. Hemoglobin A1C moderately influences the distributions and follow-up has the least effect on the distribution. These findings are consistent with the existing literature on Type 2 diabetes mellitus and more convincing than the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates.
The contradictory conclusions from the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and ${\mathcal{S}}_{Y|{\bf X}}$ estimates highlights the need to look beyond a specific location in the distribution. In the original study, patients with basal glycaemia $\geq 126mg/dL$ and/or haemoglobin A1c $\geq 6.5\%$, were diagnosed as diabetics. Therefore, to gain more insights, we plotted the predicted distributions based on the sufficient predictors $\hat{\beta}^{\top}{\bf X}$ for the different metrics. Where the Wasserstein estimates are similar, we choose only one of the two.
By examining the estimated distributions in Figure 2, we see that the distributions of patients diagnosed with diabetes do not necessarily standout. The peaks of the distributions as well as the right tails appear to vary more with the sufficient predictor for the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ based on the Wasserstein metric than the one based on the Hellinger distance. This buttresses the point that metric choice plays a crucial role, especially when dealing with the mean space, as we observed in the numerical studies. We observe a similar pattern in the predicted distributions based on the ${\mathcal{S}}_{Y|{\bf X}}$ sufficient predictors in Figure 6.

### v2

@@ Abstract
Fréchet regression is becoming a mainstay in modern data analysis for analyzing non-traditional data types belonging to general metric spaces. This novel regression method is especially useful in the analysis of complex health data such as continuous monitoring and imaging data. Fréchet regression utilizes the pairwise distances between the random objects, which makes the choice of metric crucial in the estimation. In this paper, existing dimension reduction methods for Fréchet regression are reviewed, and the effect of metric choice on the estimation of the dimension reduction subspace is explored for the regression between random responses and Euclidean predictors. Extensive numerical studies illustrate how different metrics affect the central and central mean space estimators. Two real applications involving analysis of brain connectivity networks of subjects with and without Parkinson’s disease and an analysis of the distributions of glycaemia based on continuous glucose monitoring data are provided, to demonstrate how metric choice can influence findings in real applications.
Keywords: $L_{p}$ spaces, brain imaging, Parkinson’s disease, diabetes mellitus, continuous glucose monitoring
@@ 1 Introduction
Because $Y$ is allowed to belong to a general metric space, including those that do not satisfy vector space properties, Fréchet regression is applicable in many fields. Some applications of this novel regression can be found in the study of mortality profiles in Petersen & Müller (2019), the influence of age on the development of brain myelination in Petersen et al. (2019), and the association between age, height, and weight on the structural connectivity of the human brain in Soale & Dong (2023), to name a few.
@@
The rest of the paper is organized as follows. In Section 2, we provide a brief background on metric and pseudometric spaces. We also discuss some popular metric spaces and the different metric choices for each space. In Section 3, we give a general background to Fréchet sufficient dimension reduction, followed by an extensive simulation studies in Section 4. We present two real applications in Section 5 and conclude the paper in Section 6 with some discussion. All other technical details are relegated to the Appendix.
@@ 2.1 General background
Furthermore, it is important to note that regardless of whether we are dealing with a metric or a pseudometric space, different metric functions in a given space can have dramatic effects on the geometry of the space. For instance, in the normed space, we use the $\ell_{p}$ norm, denoted as .$\lVert.\rVert_{p}$, where $p\geq 1$ is some real number. In this space, the choice of $p$ determines the geometry as shown in Figure 1. We also see from Figure 1, some “nesting” or “ordering” behavior of the $\ell_{p}$ norm.
@@
2.2 Popular metric functions in some common metric spaces
2.2.1 Euclidean space
In this space, the response set $\bm{\Omega}_{Y}\in\mathbb{R}^{r}$, where $r\geq 1$ is some positive integer. The most common metric in this space is the $\ell_{p}$ norm, which as shown in Figure 1 exhibits nesting.
Proposition 1.
@@ 2.2.1 Euclidean space
By Proposition 1, $\ell_{2}$ is the most restrictive of the $\ell_{p}$ norms, and in a finite normed space, the richness of $\mathbb{L}_{p}$ depends on the choice of metric. In particular, for $1\leq p_{1}<p_{2}\leq 2,\ \mathbb{L}_{p_{2}}\subset\mathbb{L}_{p_{1}}$ [Matoušek (2013)]. This containment relation can have significant effects on the central space estimation as we will show later. This proposition also suggests that metrics in different metric spaces that depend on the $\ell_{p}$ norm will exhibit some of this behavior.
Here, we will focus on the $\ell_{1}$ and $\ell_{2}$ norms. Thus, by Proposition 1, in the finite Euclidean space, $\mathbb{L}_{2}(Y)\subset\mathbb{L}_{1}(Y)$. Besides these two metrics, for $r\geq 2$, we will also consider the Mahalanobis distance, which is a pseudometric that takes into account the dependence between the responses. The Mahalanobis distance is given by
@@
2.2.2 Space of probability distributions
@@
By Proposition 3, when there is 1st-order stochastic dominance, $W_{1}(.)$ simply reduces to the difference in means. For example, for any two log-normal distributions $F_{1}(\mu_{1},\sigma)$ and $F_{2}(\mu_{2},\sigma)$, $F_{1}$ dominates $F_{2}$ if $\mu_{1}>\mu_{2}$. In such scenarios, using the $W_{1}(.)$ is preferable to $W_{2}(.)$ as the latter will likely introduce some noise in finite samples based on the decomposition in (4).
@@
2.2.3 Space of symmetric positive definite matrices
@@ 2.2.4 Space of networks
$D_{d}(G_{1},G_{2})=\max_{\uptau}\big{(}\lVert\mathrm{expm}(-\uptau\L_{1})-\mathrm{expm}(-\uptau\L_{2})\rVert_{F}^{2}\big{)}^{1/2},$ (7)
where Ł$\L_{i}$ is the Laplacian of graph $G_{i}$, $\uptau\in(0,1)$, and $\mathrm{expm}({\bf A})$ is the exponential of matrix ${\bf A}$ given by $\mathrm{expm}({\bf A})=\displaystyle\sum_{n=0}^{\infty}\cfrac{{\bf A}^{n}}{n!}$. Unlike the degree centrality distance, the diffusion distance can account for edge weight and other network characteristics beside node characteristics.
@@ 3.1 A brief background
Definition 2 (Metric embedding).
@@
Theorem 1 emphasizes the importance of choosing the appropriate metric as it relates to information about the central space such as dimensionality, unbiasedness, Fisher consistency, and exhaustiveness. See Li (2018) for more details on these properties.
Theorem 2.
@@
For Theorem 2 to be satisfied, it suffices for the response embedding to be Lipschitz, which may not hold for all metrics defined on the same space. For instance, the Wasserstein distance between two normal distributions is Lipschitz but the Hellinger distance is not. Some metrics may also satisfy the Lipschitz condition locally. Moreover, for metrics without upper bounds, the embedding on scaled inputs is more likely to concentrate around the mean than the embedding based on the actual values. For instance, we expect the Frobenius norm of the difference between logged positive definite matrices to concentrate around the mean faster than the Frobenius norm between the actual matrices. Thus, it is important to evaluate metrics on case-by-case basis. The central mean space estimators are especially powerful when the concentration of the metric is Gaussian or approximately Gaussian.
@@ 4.1 Euclidean responses
Model $(n,p)$ Metric FOLS sa-OLS FSIR sa-SIR
@@ 4.2 Space of univariate distributions
Here, all three models III, IV, and V are single-index with $\eta=\beta_{1}$. From Table LABEL:tab:Dist_sims, the Hellinger distance shows significant improvement in the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates for both models III and V over the Wasserstein metrics. For model III, the Wasserstein distances are not globally Lipschitz but may be locally Lipschitz when $\beta_{1}^{\top}{\bf X}<0$. Thus, we do not expect $W_{1}$ and $W_{2}$ to concentrate around the mean by Theorem 2. On the other hand, the Hellinger distance concentrates around the center. Moreover, the Hellinger distance is better suited for capturing the tail variations compared to the Wasserstein metrics. This is one of the reasons for the difference we see for the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates for model III. However, we see less difference in the ${\mathcal{S}}_{Y|{\bf X}}$ estimates as the central space is not restricted to the mean.
In model IV, the Hellinger distance performs poorly compared to Wasserstein metrics in both the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and ${\mathcal{S}}_{Y|{\bf X}}$ estimates. The 1-Wassertein performs slightly better than the 2-Wasserstein because any pair of log-normal distributions with same scale parameters exhibit stochastic dominance. Thus, based on Proposition 3, $W_{1}(.)$ captures just the difference in the means while $W_{2}(.)$ may introduce some noise from the difference in sample scale parameters and shape. The log-normal distributions are more symmetric, which makes the Wasserstein a better metric for capturing location differences than the Hellinger distance.
@@
Model $(n,p)$ Metric FOLS sa-OLS FSIR sa-SIR
@@ 4.3 Space of positive definite matrices
Model $(n,p)$ Metric FOLS sa-OLS FSIR sa-SIR
@@ 4.4 Space of networks
Model $(n,p)$ Metric FOLS sa-OLS FSIR sa-SIR
@@
5 Real Applications
In this section, we present two applications with responses belonging to two different metric spaces. In the first application we focus on inference and in the second we focus on prediction.
5.1 Functional brain connectivity network data
Recent advances in functional magnetic resonance imaging (fMRI), has allowed us to capture in-depth structural and functional activities of the human brain. By leveraging such imaging data, we can better understand certain neurological diseases such as Alzheimer’s and Parkinson’s disease (PD). In this study, we analyze the correlations between different regions of the brain generated using the HarvardOxford Parcellation method. The analysis data consists of brain connectivity networks in the Neurocon dataset released by Badea et al. (2017). There are 41 subjects in the dataset, 26 of whom have Parkinson’s disease. Details about the preprocessing of the images as well as the codes are publicly available at https://github.com/brainnetuoa/data_driven_network_neuroscience and https://doi.org/10.17608/k6.auckland.21397377.
Our response of interest is the correlation between the 48 regions of interest (ROI) of the subject’s brain. Figure 2 shows the comparison between the correlation plot of a subject with and without Parkinson’s disease.
Figure 2: Correlation plot between the 48 different regions of the brain for a subject with and without Parkinson’s disease
We take as predictors, the recorded age, sex (F/M), and diagnosis (PD/control). The age range is between 45 and 86, thus, to avoid undue influence from outliers, we use the natural logarithm of the age variable in the analysis. Next, we proceed to find the central mean space ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates using the the Fréchet OLS (FOLS) and surrogate-assisted OLS (sa-OLS) under different metrics, i.e., the Frobenius (F), Cholesky Frobenius (Chol F), and log Frobenius (log F) metrics. In particular, we want to compare the average difference in correlation structure between the subjects diagnosed with Parkinson’s disease vs those without the disease after controlling for age and sex. The results are provided in Table 5.
Table 5: ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimates based FOLS and sa-OLS under different metrics. The bootstrap standard errors and confidence intervals within two standard deviations are given in parentheses.
Predictor FOLS (F) sa-OLS (F) FOLS (Chol F) sa-OLS (Chol F) FOLS (log F) sa-OLS (log F)
Diagnosis [PD] 0.6360 (0.3842) 0.6403 (0.0456) 0.8153 (0.3654) 0.7979 (0.0427) 0.8015 (0.3708) 0.7907 (0.0460)
(-0.1324, 1.4044) (0.5491, 0.7315) (0.0845, 1.5461) (0.7125, 0.8833) (0.0599, 1.5431) (0.6987, 0.8827)
$\log$ Age -5.7125 (0.8634) -5.7132 (0.0419) -5.7654 (0.6876) -5.7660 (0.0404) -5.7636 (0.9290) -5.7626 (0.0406)
(-7.4393, -3.9857) (-5.7970, -5.6294) (-7.1406, -4.3902) (-5.8468, -5.6852) (-7.6216, -3.9056) (-5.8438, -5.6814)
Sex [F] 1.0446 (0.3873) 1.0437 (0.0016) 0.9286 (0.3633) 0.9378 (0.0032) 0.9396 (0.3708) 0.9469 (0.0050)
(0.2700, 1.8192) (1.0405 , 1.0469) (0.2020, 1.6552) (0.9314, 0.9442) (0.1980, 1.6812) (0.9369, 0.9569)
In Table 5, we see that based on the FOLS estimate with the Frobenius norm, there is no significant difference in the correlation structure between the subjects diagnosed with Parkinson’s disease and those without the diseases at a fixed age and sex. However, all other methods and metrics show significant difference. There is also variation in the level of the average difference depending on the metric choice. While both age and sex appear to be significantly related to the brain connectivity, the level of association also varies depending on the method and metric choice. Overall, the surrogate-assisted estimates appear to be more stable across metrics with smaller standard errors for the estimates compared to the Fréchet OLS estimates. The results in Table 5 emphasizes the need to explore different estimators and metrics in applications.
5.2 Glucose monitoring data
@@
Type 2 diabetes mellitus (T2DM) disease has become rampant across the world in recent decades. Classifying a patient as diabetic is typically based on some threshold of the level of glucose in the blood (glycaemia) or hemoglobin A1C, which could potentially lead to incorrect diagnosis. However, with recent advances in continuous glucose monitoring systems, we could improve the metrics for diagnosing type 2 diabetes mellitus. The goal of this study is to analyze the distribution of glycaemia and determine the factors influencing the distribution. This way, we take into account all aspects of the distribution including the center, the tails, and everything in-between.
@@
Our data comes from the continuous glucose monitoring of 208 patients selected from the University Hospital of Móstoles, in Madrid, during an outpatient visit from January 2012 to May 2015. The data consists of the distribution of glycaemia recorded every 5 seconds for each patient over at least a 24 hour period. Patients were followed until either the diagnosis of T2DM or end of the study. Patients with basal glycaemia $\geq 126mg/dL$ and/or haemoglobin A1c $\geq 6.5\%$, were diagnosed with T2DM at the end of the study. This data is publicly available and the details about the collection procedure can be found in Colás et al. (2019).
For our analysis, we take the glycaemia distributions as the response and clinical variables including age (years), body mass index (BMI; $kg/m^{2}$), basal glycaemia ($mg/dL$), and haemoglobin A1c (HbA1c; %) as the predictors. Next, we proceed to estimate the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ using the Fréchet OLS (FOLS) and surrogate-assisted OLS (sa-OLS); and the ${\mathcal{S}}_{Y|{\bf X}}$ using the Fréchet SIR (FSIR) and surrogate-assisted SIR (sa-SIR) under different metrics, i.e., the 1-Wasserstein ($W_{1}$), 2-Wasserstein ($W_{2}$), and Hellinger ($H$) distances.
@@
To compare the performance of the methods and metrics, we estimate the leave-one-out average prediction error for the conditional Fréchet mean based on the estimated sufficient predictor using model (1). Let $\hat{\beta}$ denote the basis estimate. The average leave-one-out prediction error is given by
@@
where $\hat{m}_{\oplus}(\bm{x})$ is the sample estimate of $m_{\oplus}(\bm{x})$ and $\bm{x}^{(-i)}$ is $\bm{x}$ without the $i$th observation. The results for ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and ${\mathcal{S}}_{Y|{\bf X}}$ estimates are provided in Tables 6 and 7, respectively.
Table 6: ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ estimate based on different metrics
@@
Age -0.5746 -0.5554 -0.6745 0.6590 -0.3691 0.4407
BMI 0.0271 0.0418 -0.1207 0.1055 -0.1427 0.4233
Basal glycaemia -0.4789 -0.4919 -0.4202 0.4371 0.9241 -0.9097
HbA1c -0.4283 -0.4346 -0.3658 0.3703 -0.5121 0.2622
$\delta$ 11.9933 11.9840 11.9600 11.9756 12.0355 12.1097
In Table 6, we see very similar estimates based on the Wasserstein metrics as opposed to the Hellinger distance. Going by the estimates based on the Wasserstein metrics, the most important factor that influence the glycaemia distribution is Age, followed by basal glycaemia, then Hemoglobin A1C, and lastly BMI. However, based on the Hellinger distances, the most important determinant is the basal glycaemia, followed by age, then BMI, and lastly the hemoglobin A1C. In terms of prediction accuracy, the Wasserstein metrics appear to be better than the Hellinger estimates. We visualize the predicted distributions for the best and worse predictions in Figure 3.
Similarly, in Table 7, we provide the ${\mathcal{S}}_{Y|{\bf X}}$ estimates. Again, we see conflicting results based on the Wasserstein metrics and the Hellinger distances. For the estimates based on the Wasserstein metrics, BMI is the most important, follow by age, then basal glycaemia, and then hemoglobin AIC. However, with the Hellinger distance, it is basal glycaemia, followed by hemoglobin A1C, then BMI, and lastly age. In terms of prediction, the 1-Wasserstein and Hellinger distances appear to perform better than the 2-Wasserstein. We visualize the surrogate-assisted SIR estimate based on the 1-Wasserstein and the Fréchet SIR estimate based on the Hellinger distance in Figure 7.
Table 7: ${\mathcal{S}}_{Y|{\bf X}}$ estimate based on different metrics
@@
Age -0.6268 -0.6237 -0.7085 -0.6730 -0.1396 0.0932
BMI -0.6546 -0.6471 -0.6767 -0.6987 0.3607 -0.4169
Basal glycaemia 0.5575 0.6133 0.4396 0.4989 0.6080 -0.6510
HbA1c -0.3406 -0.2906 -0.2703 -0.2442 0.5409 -0.4496
$\delta$ 11.9795 11.9567 12.0099 12.0451 11.9848 11.9909
Notice that variable importance varies depends on the estimator. The OLS estimators, i.e., FOLS and sa-OLS focus on the mean of the distribution, while the SIR estimators, i.e., FSIR and sa-SIR take the entire distributions into account. Thus, for the OLS estimators the tail variations in the distributions are not important, which is why the Hellinger distance performed worse at the predictions. On the other hand, SIR considers all aspects of the distributions, which is why the Hellinger distance becomes competitive and even performs better than the 2-Wasserstein. Another thing to notice is that estimators with significantly different basis estimates can yield similar prediction accuracy as seen in the 1-Wasserstein and Hellinger SIR estimates. Finally, by examining the estimated distributions, we see that the distributions of patients diagnosed with T2DM do not necessarily standout compared to those without the disease.